# Explanation User Interfaces: A Systematic Literature Review

### v1

@@ 2.4. Existing Literature Reviews
Many Systematic Literature Reviews (SLRs) and surveys have explored topics closely related to our SLR, underscoring the growing relevance and urgency of advancing XAI to ensure its practical implementation in real-world scenarios. These reviews often adopt distinct perspectives, focusing on different aspects of the topic. Several studies emphasise the concept of interactivity: (Chromik and Butz, 2021) focuses on interaction design principles for XUIs, offering a structured analysis of user interactions and design guidelines. Similarly, (Bertrand et al., 2023b) categorises interactive explanation techniques based on user evaluation metrics, focusing on interactivity and its effects on user perception. (Raees et al., 2024) further advances the argument by examining interactivity in XAI systems through the lens of Human-Centred AI literature, proposing the evolution of XAI into Interactive AI. Another study (Muralidhar et al., 2023), examines the components of transparency in AI systems, aiming to enhance user interpretability and proposing strategies for designing systems that are more interpretable and explainable to users. The evaluation of XAI applications through user studies represents another important area of focus. (Rong et al., 2024) investigates how these studies are conducted, while (Al-Ansari et al., 2024) analyses the role of HCI techniques in improving XAI goals. HCI remains central to the exploration of explanation design in (Ferreira and Monteiro, 2020), whose main purpose is to investigate who the recipients of AI explanations are, the motivations for providing them, and the methods for aligning explanations with user needs. The intersection between XAI and HCI is also investigated in (Alaqsam and Sas, 2024), where the authors highlight the opportunities for HCI concerning XAI tools. Furthermore, studies like those by (Laato et al., 2022) and (Haque et al., 2022)focus on how explanations can be effectively tailored to specific user categories, particularly non-technical and end users, addressing critical gaps in accessibility and usability.
@@
Finally, (de Souza Filho et al., 2024) narrows its scope to reliance-aware explainable user interfaces, synthesising secondary studies to highlight design solutions to foster appropriate reliance in XAI-assisted decision-making. Our review examines the literature on XUI, leveraging both algorithmic foundations and human-centred design principles. Specifically, our research investigates three aspects: (i) the contextual elements that guide the design of XUI solutions, (ii) AI and XAI methodologies commonly employed in the development of XUIs, and (iii) the evaluation methods, driven by HCI principles, used to assess XUIs effectively. Moreover, we also discuss the design principles that emerge from the literature. Unlike prior works that often treat these aspects in isolation, our review seeks to integrate these dimensions to identify design trends among human-centred XUIs and XAI algorithms by synthesising insights from 146 papers retrieved from IEEE, ACM, and Scopus. Our contribution lies in providing a comprehensive overview that progresses the discussion on XUIs by connecting the technical, design, and evaluation components in a unified way, addressing the need for integrated solutions in real-world XAI applications.
@@ 3. Planning and Conducting the Systematic Literature Review
We used a rigorous, reproducible methodology to perform an SLR. According to Kitchenham, conducting an SLR involves three main stages: planning, execution, and reporting (Kitchenham, 2004). This section details the first two stages, while the third one is covered in Sections 4, 5, 6 and 7.
@@ 3.1. Planning the SLR
Planning an SLR is a multi-step process that involves the following activities (Kitchenham, 2004):
  1. (1)
    formulating a research question;
  2. (2)
    defining the set of search strings;
  3. (3)
    electing the data sources;
  4. (4)
    defining the inclusion and exclusion criteria.
In this section, we report the details of each activity, discussing the motivations behind every choice.
3.1.1. Formulation of the Research Question
@@
    What XAI solutions are used to develop XUIs?
@@ 3.1.2. Definition of the Search Strings
This formulation ensures a comprehensive yet precise retrieval of relevant studies, aligning with the scope of our SLR.
@@ 3.1.3. Selection of Data Sources
We started this SLR from scientific digital libraries. The examined libraries were ACM Digital Library [^1] https://dl.acm.org, IEEE Xplore [^2] https://ieeexplore.ieee.org, and Scopus [^3] https://www.scopus.com/. During the definition of the search string, we noticed that the same search has to be performed differently depending on the library (i.e., using a different syntax). Each library contained various options for searching content: for example, they allow searching for keywords within the article title, abstract, full text, or all of the above. We searched for the most comprehensive choice available for each digital library. All searches were conducted on the entire database due to the inherent multidisciplinarity of XUIs applications.
@@ 3.1.4. Definition of the Inclusion and Exclusion Criteria
This step concerns the final selection of the relevant publications. Kitchenham distinguishes between practical inclusion and exclusion criteria, which address logistical constraints, and quality-based criteria, which are derived from the research question to refine the selection of relevant studies. Table 1 outlines the practical criteria applied in this study. Notably, we focused on publications starting from 2013 since XAI was formalized only in 2017 (Gunning and Aha, 2019), but concerns on biases and the need for explanations started earlier, as discussed in (Minsky et al., 2013).
Additionally, inclusion criteria were established to ensure the selected papers aligned with the research objectives. These criteria focus on the content of the research question and require that a study:
  • Involves a user study as part of its evaluation;
  • Incorporates an explanation algorithm;
  • Is based on an AI model, or, if the AI system is not explicitly defined, is designed with AI in mind (e.g., Wizard of Oz studies).
@@ 3.2.1. Literature Review Execution
The initial search across the digital library yielded a total of 7084 potentially relevant publications. After a duplicate check, a dataset of 6095 publications was obtained. After filtering, using the previously defined inclusion criteria, 1023 publications were excluded as they did not meet the quality criteria, leaving a final dataset of 5072 candidates. After this first filtering step, the title and abstract were analyzed, resulting in the exclusion of a further 4658 publications, reducing the number to 414. Filtering was conducted independently by three researchers, and conflicts were treated as relevant and moved to the next step. The publications were then carefully analyzed by accessing their full text, narrowing down to 142 valid publications for the review. After forward and backward snowballing, an additional 148 candidates were retrieved and filtered down to 4 using the same inclusion criteria. In conclusion, the final dataset comprises a total of 146 publications.
@@ 3.2.2. Data Synthesis
For each paper, multiple categorical dimensions were collected for analysis, which align with established taxonomies to facilitate a structured evaluation of the results. A complete description of the dimensions and their corresponding research questions is presented in table 2.
@@
Dimension Research Questions Description
Application domain RQ1 Refers to the specific field or sector where the Explainable AI (XAI) system is applied, such as healthcare, finance, or education.
Type of data RQ1 Indicates the nature of the data used by the AI system, such as tabular, image, text, or audio.
AI RQ2 The specific AI model or architecture employed, such as neural networks, decision trees, or ensemble methods.
AI Task RQ2 Describes the goal of the AI system, such as classification, regression, clustering, or prediction.
Output Type RQ2 Refers to the format of the AI model’s outputs, such as probabilities, labels, and rankings.
XAI Techniques RQ2 The methods used to generate explanations, such as counterfactual reasoning, feature importance, and Shapley values.
Explanation Modality RQ2 The mode of explanation delivery, such as textual, visual, natural language, or a combination of these.
Interactivity RQ2 Indicates whether the XUI system implements interactive elements to enhance user engagement and understanding.
Visual Mode RQ2 The type of visual representation used in the interface, such as heatmaps, trendlines, or other visual models.
Type of Study RQ3 The methodological approach used to evaluate the Explanation User Interface, such as interviews, controlled experiments, or user observations.
Type of Users RQ1, RQ3 The target audience for the system, such as domain experts, end-users, or AI specialists.
Number of Participants RQ3 The sample size of users involved in studies evaluating the system.
Metrics RQ3 The criteria used to assess the effectiveness of the XAI system, such as trust, usability, understanding, or task performance.
The knowledge extracted from the selected literature was analyzed in the following four sections (sections 4, 5, 6 and 7), one per research question.
4. RQ1. What influences the design of XUIs?
This research question investigates what influences XUIs design, identifying the critical elements (e.g., constraints and considerations) that guide it. Our analysis examines the problem space from two perspectives. From a human-centred perspective, we focus on the domain of the application and the target users to whom the interface is designed. From an algorithmic perspective, we analyse the underlying AI layer, focusing on the type of data utilised.
Table 3. Distribution of User Types in the Reviewed Literature
@@
Table 4. Distribution of Application Domains
@@
Natural Language Processing 13
Artificial Intelligence and Robotics Systems 12
Media and Communication 10
@@
Agriculture and Nature 2
Table 5. Distribution of Data Types
@@
4.1. Application Domains
The analysis of application domains, reported in Table 5, highlights the prominence of XUIs in high-stakes fields such as health and finance (He et al., 2024; Esfahani et al., 2024; Barda et al., 2020; Wang et al., 2021; Kim et al., 2023b; Bhattacharya et al., 2024; Meza Martínez et al., 2023; Purificato et al., 2023). These domains, characterised by their critical decision-making requirements, align closely with the ones in which XAI is extremely useful (Combi et al., 2022). However, underexplored areas like mobility (Antar et al., 2024; Schneider et al., 2021b), cybersecurity (Desolda et al., 2023; Stites et al., 2021; Wu et al., 2021), and education (Shen et al., 2023; Shin et al., 2022) suggest opportunities for expanding the impact of XUIs into new contexts. The number of domains reinforces the broad applicability of XUIs solutions across various sectors.
4.2. Type of Users
@@
4.3. Data Types
The distribution of data types in XUI designs is presented in 5. Tabular and image data prevail, reflecting the prevalence of structured datasets and the importance of images in the most-explored domain (i.e., medicine) (Hwang et al., 2022a; Gromowski et al., 2020). This also highlights the importance of visual representation in explainability. Other types, like text (Riveiro and Thill, 2021) and time series (Upasane et al., 2021), indicate a growing interest in different data modalities. However, the relatively smaller focus on multimodal and media-based data (Wang et al., 2022) reveals potential for future exploration, particularly in complex scenarios requiring the integration of multiple data types.
4.4. Interplay Among Application Domain, User Type, and Data Type
@@
Heatmap showing the relationship between user types and application domains, illustrating how different user categories are distributed across various domains
Similarly, the finance and economics domain demonstrates a dual focus, catering to both domain experts (Purificato et al., 2023; Füßl et al., 2024), such as financial analysts, and non-experts, such as general users of financial services (Bove et al., 2022; Chromik, 2021; Bertrand et al., 2023a). Most of the time, XUI systems in this domain focus on time series and tabular data. This reflects the need for XUIs that provides clear, interpretable insights for specialists while ensuring accessibility and usability for broader audiences.
Media and communication domain displays a balanced representation of general users and domain experts (without a strong focus on a specific data type) (Anderson et al., 2020; Wu et al., 2022; Khanna et al., 2022), underscoring the necessity of designing interfaces that make AI-driven explanations accessible to diverse audiences, including professionals and general users.
@@
However, AI and robotic systems are predominantly associated with AI and domain experts (Piorkowski et al., 2023; Baniecki et al., 2024; Nourani et al., 2022a; Mishra et al., 2022). This trend aligns with the technical expertise required to develop, understand, and evaluate such systems. For general-purpose applications, the findings indicate a noticeable emphasis on AI experts, suggesting that such interfaces often prioritise technical depth and flexibility. These systems are frequently foundational tools or frameworks designed to be adaptable across various domains, appealing to users with advanced AI knowledge who can tailor them to specific use cases.
@@
5. RQ2. What XAI solutions are used to develop XUIs?
@@
This research question examines the connection between the algorithms used to generate explanations and how these are conveyed through XUIs. To structure the analysis, we utilised the taxonomy proposed by Guidotti et al. (Guidotti et al., 2019), as outlined in Section 2.1, which categorises XAI techniques based on their methodological characteristics and their application to AI models. Additionally, we consider the framework described by Gunning and Aha to classify explanations modalities (i.e., how explanations are conveyed to users through different mediums) (Gunning and Aha, 2019).
5.1. Relationships Between AI Systems and Explainability Techniques
Table 7 summarises the AI algorithms employed in the surveyed studies. Neural networks prevail over the other techniques, followed by ensemble methods and white-box models. These results highlight the prevalence of complex, opaque models like neural networks, which require XAI techniques to enhance interpretability.

### v2

@@ 2.4. Existing Literature Reviews
Many Systematic Literature Reviews (SLRs) and surveys have explored topics closely related to our SLR, underscoring the growing relevance and urgency of advancing XAI to ensure its practical implementation in real-world scenarios. These reviews often adopt distinct perspectives, focusing on different aspects of the topic. Several studies emphasise the concept of interactivity: (Chromik and Butz, 2021) focuses on interaction design principles for XUIs, offering a structured analysis of user interactions and design guidelines. Similarly, (Bertrand et al., 2023b) categorises interactive explanation techniques based on user evaluation constructs, focusing on interactivity and its effects on user perception. (Raees et al., 2024) further advances the argument by examining interactivity in XAI systems through the lens of Human-Centred AI literature, proposing the evolution of XAI into Interactive AI. Another study (Muralidhar et al., 2023) examines the components of transparency in AI systems, aiming to enhance user interpretability and proposing strategies for designing systems that are more interpretable and explainable to users. The evaluation of XAI applications through user studies represents another important area of focus. (Rong et al., 2024) investigates how these studies are conducted, while (Al-Ansari et al., 2024) analyses the role of HCI techniques in improving XAI goals. HCI remains central to the exploration of explanation design in (Ferreira and Monteiro, 2020), whose main purpose is to investigate who the recipients of AI explanations are, the motivations for providing them, and the methods for aligning explanations with user needs. The intersection between XAI and HCI is also investigated in (Alaqsam and Sas, 2024), where the authors highlight the opportunities for HCI concerning XAI tools. Furthermore, studies like those by (Laato et al., 2022) and (Haque et al., 2022)focus on how explanations can be effectively tailored to specific user categories, particularly non-technical and end users, addressing critical gaps in accessibility and usability. Finally, (de Souza Filho et al., 2024) narrows its scope to reliance-aware explainable user interfaces, synthesising secondary studies to highlight design solutions to foster appropriate reliance in XAI-assisted decision-making.
Novelty of our SLR
Unlike other research, our review examines the literature on XUI, leveraging both algorithmic foundations and human-centered design principles. Specifically, the current research investigates three aspects: i. the contextual elements that guide the design of XUI solutions; ii. AIand XAI methodologies commonly employed in the development of XUIs; iii. the evaluation methods, driven by HCI principles, used to assess XUIs effectively.
@@
Moreover, we also discuss the design principles that emerge from the literature. Unlike prior works that often treat these aspects in isolation, our review seeks to integrate these dimensions to identify design trends among human-centred XUIs and XAI algorithms by synthesising insights from 146 papers retrieved from IEEE, ACM, and Scopus. Our contribution lies in providing a comprehensive overview that progresses the discussion on XUIs by connecting the technical, design, and evaluation components in a unified way, addressing the need for integrated solutions in real-world XAI applications.
@@ 3. Planning and Conducting the Systematic Literature Review
We used a rigorous, reproducible methodology to perform an SLR. According to Kitchenham, conducting an SLR involves three main stages: planning, execution, and reporting (Kitchenham, 2004). This section details the first two stages, while the third one is covered in Sections 4.1, 4.2, 4.3 and 4.4.
@@ 3.1. Planning the SLR
Planning an SLR is a multi-step process that involves the following activities (Kitchenham, 2004): (1) formulating a research question; (2) defining the set of search strings; (3) electing the data sources; (4) defining the inclusion and exclusion criteria. In this section, we report the details of each activity, discussing the motivations behind every choice.
3.1.1. Formulation of the Research Questions
@@
    What XAI solutions (e.g., tools, frameworks, AI models) are used to develop XUIs?
@@ 3.1.2. Definition of the Search Strings
This formulation ensures a comprehensive yet precise retrieval of relevant studies, aligning with the scope of our SLR. In particular, our goal was to identify papers that simultaneously address all three areas — eXplainable Artificial Intelligence, Human-Centered Design, and Interactive User Interfaces — rather than focusing on each aspect in isolation. These three dimensions are further operationalized and analyzed according to the contribution types described in Section 3.2.2.
@@ 3.1.3. Selection of Data Sources
We started this SLR from scientific digital libraries. The examined libraries were ACM Digital Library [^1] https://dl.acm.org, IEEE Xplore [^2] https://ieeexplore.ieee.org, and Scopus [^3] https://www.scopus.com/. During the definition of the search string, we noticed that the same search has to be performed differently depending on the library (i.e., using a different syntax). Each library contained various options for searching content: for example, they allow searching for keywords within the article title, abstract, full text, or all of the above. We searched for the most comprehensive choice available for each digital library. All searches were conducted on the entire database due to the inherent multidisciplinarity of XUIs applications. Google Scholar was omitted, as it does not meet the inclusion criteria based on peer-reviewed publications and venue quality. Moreover, since the selected databases already cover most high-quality venues indexed by Google Scholar, its inclusion would mainly have increased the number of duplicates without improving coverage.
@@ 3.1.4. Definition of the Inclusion and Exclusion Criteria
This step concerns the final selection of the relevant publications. Kitchenham distinguishes between practical inclusion and exclusion criteria, which address logistical constraints, and quality-based criteria, which are derived from the research question to refine the selection of relevant studies. Table 1 outlines the practical criteria applied in this study. Notably, we focused on publications starting from 2013 since XAI was formalized only in 2017 (Gunning and Aha, 2019), but concerns on biases and the need for explanations started earlier, as discussed in (Minsky et al., 2013). Additionally, inclusion criteria were established to ensure the selected papers aligned with the research objectives. These criteria focus on the content of the research question and require that a study:
(i) Involves a user study as part of its evaluation; (ii) Incorporates an explanation algorithm; (iii) Is based on an AI model, or, if the AI system is not explicitly defined, is designed with AI in mind (e.g., Wizard of Oz studies).
@@ 3.2.1. Literature Review Execution
The initial search across the digital library yielded a total of 7084 potentially relevant publications. After a duplicate check, a dataset of 6095 publications was obtained. After filtering, using the previously defined inclusion criteria, 1023 publications were excluded as they did not meet the quality criteria, leaving a final dataset of 5072 candidates. After this first filtering step, the title and abstract were analyzed, resulting in the exclusion of a further 4658 publications, reducing the number to 414. Filtering was conducted independently by three researchers; any conflict was reported and addressed in the next step. Disagreements were resolved through discussion among the researchers, and in cases where one researcher disagreed with the other two, a majority vote was used. Inter-rater agreement on inclusion/exclusion decisions was considerable, with a Cohen’s $\kappa$ of .81.81. Through successive discussions, full inter-rater agreement was reached. The publications were then carefully analyzed by accessing their full text, narrowing down to 142 valid publications for the review. After forward and backward snowballing, an additional 148 candidates were retrieved and filtered down to 4 using the same inclusion criteria. In conclusion, the final dataset comprises a total of 146 publications.
@@ 3.2.2. Data Synthesis
For each paper, multiple categorical dimensions were collected for analysis, which align with established taxonomies to facilitate a structured evaluation of the results. Furthermore, for papers that reported insights for designing XUI systems—such as principles, recommendations, lessons learned, etc.—we collected design guidelines and frameworks. The knowledge extracted from the selected literature is reported in Section 4.
4. Results
The 13 dimensions identified in the literature review predominantly reflect either a human-centered or algorithmic perspective. For example, aspects related to user validation (e.g., visual mode, type of study) belong to the former category, while aspects related to the AI model (e.g., AI task, type of data) fall into the latter. The only exception is the evaluation construct dimension, which is equally applicable to both perspectives. Moreover, data from each dimension were used to answer one or more research questions. A complete description of the extracted dimensions and their corresponding research questions is presented in Table 2.
@@ 4. Results
Dimension Perspective Research Questions Description
Application domain Human-centered RQ1 Refers to the specific field or sector where the Explainable AI (XAI) system is applied, such as healthcare, finance, or education.
Type of data Algorithmic RQ1 Indicates the nature of the data used by the AI system, such as tabular, image, text, or audio.
AI Algorithmic RQ2 The specific AI model or architecture employed, such as neural networks, decision trees, or ensemble methods.
AI Task Algorithmic RQ2 Describes the goal of the AI system, such as classification, regression, clustering, or prediction.
Output Type Algorithmic RQ2 Refers to the format of the AI model’s outputs, such as probabilities, labels, and rankings.
XAI Techniques Algorithmic RQ2 The methods used to generate explanations, such as counterfactual reasoning, feature importance, and Shapley values.
Explanation Modality Human-centered RQ2 The mode of explanation delivery, such as textual, visual, natural language, or a combination of these.
Interactivity Human-centered RQ2 Indicates whether the XUI system implements interactive elements to enhance user engagement and understanding.
Visual Mode Human-centered RQ2 The type of visual representation used in the interface, such as heatmaps, trendlines, or other visual models.
Type of Study Human-centered RQ3 The methodological approach used to evaluate the Explanation User Interface, such as interviews, controlled experiments, or user observations.
Type of Users Human-centered RQ1, RQ3 The target audience for the system, such as domain experts, end-users, or AI specialists.
Number of Participants Human-centered RQ3 The sample size of users involved in studies evaluating the system.
Evaluation constructs Both RQ3 The criteria used to assess the effectiveness of the XAI system, such as trust, usability, understanding, or task performance.
4.1. RQ1. What influences the design of XUIs?
This research question investigates the factors that influence XUIs design, identifying the critical elements (e.g., constraints and considerations) that guide it. Our analysis examines the problem space from two perspectives. From a human-centred perspective, we focus on the domain of the application and the target users to whom the interface is designed. From an algorithmic perspective, we analyse the type of data utilised by the underlying AI layer.
Summary of findings
High-stakes domains (e.g., health, finance) more commonly employ XUIs, which are generally used by domain experts (e.g., physicians). Moreover, AI systems behind XUIs typically work with tabular and image data (e.g., patient data, medical scans). XUIs used in the finance domain, albeit high-stakes, are also intended for non-expert users and tend to utilize time-series and tabular data.
Table 3. User Types Dist.
@@
Table 4. App. Domains Dist.
@@
NLP 13
AI & Robotics 12
Media & Comm. 10
@@
Agriculture 2
Table 5. Data Types Dist.
@@
4.1.1. Application Domains
The analysis of application domains, reported in Table 5, highlights the prominence of XUIs in high-stakes fields such as health and finance (He et al., 2024; Esfahani et al., 2024; Barda et al., 2020; Wang et al., 2021; Kim et al., 2023b; Bhattacharya et al., 2024; Meza Martínez et al., 2023; Purificato et al., 2023). These domains, characterised by their critical decision-making requirements, align closely with the ones in which XAI is extremely useful (Combi et al., 2022). However, underexplored areas like mobility (Antar et al., 2024; Schneider et al., 2021a), cybersecurity (Desolda et al., 2023; Stites et al., 2021; Wu et al., 2021), and education (Shen et al., 2023; Shin et al., 2022) suggest opportunities for expanding the impact of XUIs into new contexts. The number of domains reinforces the broad applicability of XUIs solutions across various sectors.
4.1.2. Type of Users
@@
4.1.3. Data Types
The distribution of data types in XUI designs is presented in Table 5. Tabular and image data prevail, reflecting the prevalence of structured datasets and the importance of images in the most-explored domain (i.e., medicine) (Hwang et al., 2022b; Gromowski et al., 2020). This also highlights the importance of visual representation in explainability. Other types, like text (Riveiro and Thill, 2021) and time series (Upasane et al., 2021), indicate a growing interest in different data modalities. However, the relatively smaller focus on multimodal and media-based data (Wang et al., 2022) reveals potential for future exploration, particularly in complex scenarios requiring the integration of multiple data types.
4.1.4. Interplay Among Application Domain, User Type, and Data Type
@@
Relationship between user types and application domains. The figure illustrates how different user categories are distributed across various domains
@@
Similarly, the finance and economics domain demonstrates a dual focus, catering to both domain experts (Purificato et al., 2023; Füßl et al., 2024), such as financial analysts, and non-experts, such as general users of financial services (Bove et al., 2022; Chromik, 2021; Bertrand et al., 2023a). Most of the time, XUI systems in this domain focus on time series and tabular data. This reflects the need for XUIs that provides clear, interpretable insights for specialists while ensuring accessibility and usability for broader audiences. The media and communication domain displays a balanced representation of general users and domain experts (without a strong focus on a specific data type) (Anderson et al., 2020; Wu et al., 2022; Khanna et al., 2022), underscoring the necessity of designing interfaces that make AI-driven explanations accessible to diverse audiences, including professionals and general users. However, AI and robotic systems are predominantly associated with AI and domain experts (Piorkowski et al., 2023; Baniecki et al., 2024; Nourani et al., 2022; Mishra et al., 2022). This trend aligns with the technical expertise required to develop, understand, and evaluate such systems. For general-purpose applications, the findings indicate a noticeable emphasis on AI experts, suggesting that such interfaces often prioritise technical depth and flexibility. These systems are frequently foundational tools or frameworks designed to be adaptable across various domains, appealing to users with advanced AI knowledge who can tailor them to specific use cases.
@@
4.2. RQ2. What XAI solutions (e.g., tools, frameworks, AI models) are used to develop XUIs?
Figure 3. Heatmap showing the comparison between AI models and XAI techniques, illustrating the most frequent usage of XAI techniques with AI models identified in the analyzed literature.
@@
This research question has a more algorithmic-centric perspective and examines the connection between the algorithms used to generate explanations and how these are conveyed through XUIs. To structure the analysis, we utilised the taxonomy proposed by Guidotti et al. (2019), as outlined in Section 2.1, which categorises XAI techniques based on their methodological characteristics and their application to AI models. Additionally, we consider the framework described by Gunning and Aha (2019) to classify explanation modalities (i.e., how explanations are conveyed to users through different media). In this context, the term solution is intentionally used in a broad sense, referring to any XUI component or approach—consistent with the DARPA XAI definition—that mediates the interaction between the user and the explanation provided by an XAI system (See Section 2.3).
Summary of findings
The most common XAI solutions in XUIs are feature importance, counterfactual explanations, and Shapley values, while the most common AI models are neural networks, ensembles, and transparent models. Some XAI solutions are more present for certain AI models (e.g., neural activation techniques are used with neural networks, explanation decision trees are used with reinforcement learning models). Most XUIs use visual-based explanations, most commonly employing heatmaps, barcharts, and trend lines. Textual explanations are also widely used. XUI systems often consist of interactive UI.
4.2.1. Relationships Between AI Systems and Explainability Techniques